OpenStudy (anonymous):

Find a second-degree polynomial P such that P(2) = 5, P'(2)=3, P''(2).

OpenStudy (anonymous):

dude let the polynomial be ax^2+bx+c find P(2) P'(2) and P"(2)

OpenStudy (anonymous):

I am not sure what that means.. I sort of need it explained...

OpenStudy (anonymous):

if P(x)=ax^2+bx+c then whats P(2)

OpenStudy (anonymous):

a2x^2+b2+c

OpenStudy (anonymous):

nope it should be a(2)^2+b(2)+c=4a+2b+c=5(given)

OpenStudy (anonymous):

2 things.. how does that = 5, and where did you get " P(x)=ax^2+bx+c"

OpenStudy (anonymous):

we assumed the equation to be that and its given that P(2)=5

OpenStudy (anonymous):

ok, but where did you get " P(x)=ax^2+bx+c" Is that some standard equation?

OpenStudy (dumbcow):

yes that is standard quadratic (2nd degree) equation

OpenStudy (anonymous):

its a general form of the quadratic equation

OpenStudy (anonymous):

and their answer to P(2)=5 was x^2-x+3

OpenStudy (anonymous):

did you find \(P'\) and \(P''\)?

OpenStudy (dumbcow):

P(x) = ax^2 +bx +c P'(x) = 2ax +b P''(x) = 2a

OpenStudy (dumbcow):

then work backwards to solve for a,b,c

OpenStudy (anonymous):

what one would i use again?

OpenStudy (anonymous):

and as per dumbcow's answer since \(P''(x)=2a\) and \(P''(x)=2\) you have \[2a=2\implies a=1\]

OpenStudy (anonymous):

wait, how does "P′′(x)=2a and P′′(x)=2 " happen?

OpenStudy (dumbcow):

wait is that P''(x) = 2 or P''(2) = ?

OpenStudy (anonymous):

what?

OpenStudy (dumbcow):

in your post what does P''(2) equal ?

OpenStudy (anonymous):

no, how is "P′′(x)=2a" and "P′′(x)=2 " the same..

OpenStudy (dumbcow):

if its given that P''(x) = 2, then you can set 2a = 2, --> a=1 P''(x) = 2a because it is 2nd derivative of P(x) do you know how to take derivative of polynomial?

OpenStudy (anonymous):

I am pretty sure i know, let me try to make sure

OpenStudy (dumbcow):

anyway there is a typo in your question.... you have P(2) =5 , P'(2) = 3, P''(2) = what??

OpenStudy (anonymous):

2

OpenStudy (dumbcow):

then satellite was correct, set 2a = 2 --> a=1

OpenStudy (anonymous):

ok

OpenStudy (dumbcow):

now set 2a(2) +b = 3 plug in a value, solve for b

OpenStudy (dumbcow):

a(2)^2 +b(2) +c = 5 plug in "a" and "b", solve for c

OpenStudy (anonymous):

ahhh, ok, now i get it!

OpenStudy (anonymous):

Thanks a million!

OpenStudy (dumbcow):

no problem